### v1

@@ Introduction
Maxwell’s equations are a four-part system of differential equations that forms the cornerstone of classical electromagnetism and describes the space-time dynamics of macroscopic electromagnetic fields. This system of equations reveals the fundamental relationships among the electric field vector E, the magnetic field vector H, the electric flux density D, the magnetic flux density B, and the current density J. The electromagnetic properties of the medium are defined by the parameters electric permittivity ($\varepsilon$), magnetic permeability ($\mu$), and electrical conductivity ($\sigma$). The differential-form expressions of the equations are as follows.
@@
In the literature, various numerical techniques such as the finite difference time domain (FDTD) [1138693], finite element method (FEM)[CLOUGH199089], and method of moments (MoM) [1447363] have been developed for modeling and analyzing electromagnetic fields. These conventional approaches are based on the discretization of Maxwell’s equations and iterative solution algorithms, and they require intensive computational resources in problems involving complex geometries and variable material properties.
In this context, Physics-Informed Neural Networks (PINNs) were introduced between 2017 and 2019 by Maziar Raissi, Paris Perdikaris, and George E. Karniadakis [RAISSI2019686]. The PINN method involves directly incorporating physical laws defined by partial differential equations into the neural network training process. Although it can be formulated within a supervised learning framework, it generally reduces dependence on labeled data since it is performed through optimization residuals.
In the literature, adaptations of PINN applications to the time-dependent Maxwell equations have remained limited due to issues such as spectral bias, the sensitivity of electromagnetic problems to scaling, causality violations, and loss balancing. The aim of this study is to present a hybrid methodology that addresses the common issues observed in PINN applications for the time-dependent Maxwell equations.
In section 1, a well-posed initial-boundary value problem formulation is obtained for the TMz mode of the time-dependent Maxwell equations. The non-dimensionalization steps are carried out and the main physics-informed loss components (PDE, boundary, and initial condition loss) used for the resulting coupled PDE system are defined.
@@
In section 3, the metrics used in the study are introduced and the performance of the PINN model is evaluated through field-based error metrics and energy metrics and the obtained results are discussed. In addition, three different variants of the Poynting regularizer developed in this study are examined under the ablation studies section.
@@ 1.1 Well-Posedness
In their most general form, Maxwell’s equations form a coupled system of partial differential equations. Without appropriate initial and boundary data, the problem is underdetermined and admits infinitely many solutions. Neural networks rely on the universal approximation theorem and can essentially be viewed as function approximators that perform function-space mapping. [HORNIK1989359]
For this reason, in order to solve Maxwell’s equations with PINN architectures, the problem that the model will learn must first be formulated in a well-posed manner. According to Hadamard’s classical approach, three criteria must be satisfied for well posedness: (1) existence of a solution, (2) uniqueness of the solution, and (3) continuous dependence of the solution on the data. [hadamard2014lectures]
In this study, the 2D time-domain Maxwell equations are considered in the $TM_{z}$ mode. A rectangular cavity with PEC walls is used, and a Gaussian pulse initialization is applied to excite the modal content. Therefore, under the source-free and lossless assumption ($J=0,\sigma=0$), and together with appropriate initial and boundary conditions, the resulting system defines a well-posed initial–boundary value problem is
@@
Modal decomposition is a common strategy in PINN formulations for electromagnetic problems, as it facilitates mode-specific learning and improves numerical conditioning. [pugin2024maxwellpinn] The process of obtaining the well-posed system of equations and the modal decomposition step are described in detail in Appendix A.
@@ 1.2 Physics-Informed Loss Components
For electromagnetic cavity resonator modeling, the loss function formulation consists of partial differential equation (PDE), initial condition (IC), and boundary condition (BC) components. The PDE loss component is defined through residual functions that enforce Maxwell’s curl laws. The boundary conditions loss, which varies depending on the specific scenario, verifies whether the equations satisfy the physical laws at the boundaries. The initial condition loss regularizes the transient electromagnetic response inside the cavity according to the chosen initialization type. In this section, the theoretical development of the loss components is addressed.
@@ 2 Implementation
However, in the implementation stage, this changes. When training PINNs, the infinite-dimensional solution space must be approximated using finite parameters and integral norms must be optimized using sampled estimators. In this discrete, parametric and local-optimization-based setting, the guarantees provided by theory may be insufficient. Scenarios that contradict physical reality are entirely plausible. For example, the optimization algorithm could violate energy conservation while minimising PDE residuals; the neural network could allow information to flow backwards in time; or field discontinuities could arise at material interfaces. These scenarios do not originate from a deficiency in the theoretical formulation, but from the intrinsic limitations of approximation and sampling-based learning in neural networks. Consequently, the main goal at the implementation stage is to translate the physical guarantees implicit in theory into explicit constraints that guide the optimizer.
@@ 2.1 Training Pipeline
The core of the model is a deep fully connected neural network with 8 hidden layers, each consisting of 128 neurons. During training, to improve gradient flow and to overcome the optimization difficulties of deep networks, a skip connection is applied after every second hidden layer inspired by the ResNet architecture. [he2015deepresiduallearningimage]
def forward(self, x: torch.Tensor, y: torch.Tensor, t: torch.Tensor):
# ... input normalization, sinusoidal time features ...
out = self.activation(self.input_layer(inp))
# Skip connections every 2 layers
for i, layer in enumerate(self.hidden_layers):
residual = out
out = self.activation(layer(out))
if i % 2 == 1:
out = out + residual
out = self.output_layer(out)
# ... output scaling for magnetic field components ...
return Ez, Hx, Hy
To ensure robust performance across different geometries, a set of problem-specific engineering strategies is adopted. The spatial coordinates are dynamically normalized to the range [1,1][-1,1] depending on the cavity aspect ratio $\beta=b/a$. To learn temporal periodic behaviors more effectively, sinusoidal feature augmentation is applied to the time input. To prevent training instabilities arising from the different magnitudes of electromagnetic field components, the magnetic field components are scaled in the output layer based on physical considerations, and the standard Xavier initialization method is used in a modified form with a lower gain factor.
@@
The optimization process is based on a two-stage strategy that combines the advantages of first-order and second-order methods. In the first stage, the global structure of the solution is learned for 1500 epochs using the Adam optimizer. In this stage, gradient clipping is used to eliminate the risk of gradient explosion, and an LR scheduler that dynamically reduces the learning rate on plateaus where convergence slows down is employed. In the second stage of optimization, starting from the point reached by Adam, high-precision fine-tuning is performed with the L-BFGS optimizer, which uses local curvature information of the loss landscape.
def train_model(model, training_data):
# Stage 1: Adam optimizer with dynamic weighting
optimizer_adam = torch.optim.Adam(model.parameters(), lr=1e-3)
scheduler = ReduceLROnPlateau(optimizer_adam, patience=100, factor=0.8)
for epoch in range(ADAM_EPOCHS):
loss = compute_total_loss(model, training_data, epoch) # epoch for dynamic weighting
loss.backward()
torch.nn.utils.clip_grad_norm_(model.parameters(), 1.0)
optimizer_adam.step()
scheduler.step(loss)
# Stage 2: L-BFGS optimizer
optimizer_lbfgs = torch.optim.LBFGS(
model.parameters(), lr=0.5,
max_iter=50, history_size=100,
line_search_fn=’strong_wolfe’
for _ in range(LBFGS_EPOCHS):
optimizer_lbfgs.step(closure)
return model
@@
To maximize both the efficiency and the accuracy of the training process, both sampling strategies and loss function weights are managed dynamically. Collocation points, at which the PDE residual is evaluated, are adapted according to the cavity aspect ratio, and their density is increased for narrow geometries.
@@
Similarly, to preserve causality during transitions between time windows, denser sampling is performed in regions close to the beginning of each window. During the early epochs of training, the weight of the loss term representing the initial conditions $(\mathcal{L}_{IC})$ is increased, while in later epochs the weight of the loss term representing the PDE $(\mathcal{L}_{PDE})$ is gradually raised. This integrated approach enables the model to first learn the overall structure of the problem and then refine its solution to achieve full consistency with the governing physical equations.
@@ 2.2 Causality of PINNs
Due to the hyperbolic nature of the Maxwell system, the field value at each point depends only on the initial and boundary data within that point’s past light cone. In other words, the field values at $t=1.0$ are formed as a function of the values at $t<1.0$. While this behavior is observed in all physical systems, it is a natural consequence of causality.
@@
Various approaches have been explored in the literature to prevent this issue. Examples include causal training [wang2022respectingcausalityneedtraining], which directly encodes causal information flow into the loss design; time-marching [Mattey_2022] and causal sweeping [Penwarden_2023], which arrange training sequentially from past to future; XPINNs-type decompositions [jagtap2020extended], which split the domain into space-time subregions; and curriculum learning [10.1145/1553374.1553380], which aligns optimization by gradually changing difficulty or sampling weights.
In this study, a time-domain decomposition strategy is adopted to prevent causality violation. This strategy is applied to the model through the time-marching approach.
@@ 2.2.1 Time Marching
The time-marching method partitions the temporal domain into sequential windows, training each window separately. This approach is inspired by the time-stepping concept in finite difference time-domain (FDTD) methods, while preserving the mesh-free advantages of PINNs.
@@ 2.2.3 Causality-Aware Weighting
One of the important issues in time-dependent PINN training is the phenomenon of forward error accumulation, where the error grows as it is propagated forward in time. When time marching is used, even if each window is solved within itself, a small PDE inconsistency at the beginning of the window can spread through derivative terms and become pronounced at mid to late times. For this reason, a causality-aware weighting mechanism is proposed in this study. This mechanism can be considered a simplified and deterministic variant of the causal training approach.
@@
In practice, with the choice $\gamma=2$, samples at the beginning of the window $\tau=0$ receive weight $1$, while those at the end $(\tau=1)$ receive approximately $e^{-2}\simeq 0.135$. This creates a weight ratio of approximately $7.4\times$ between the beginning and the end of the window. As a result, stochastic gradient estimates systematically focus on the beginning of the window, optimization suppresses early-time PDE inconsistencies first, and an equation-consistent initial regime is established.
@@ 2.3.1 Interface Continuity Loss
In more detail, when successive time windows are separated as $\mathcal{W}_{i}=[t_{i},t_{i+1}]$, the discontinuity of the fields is explicitly penalized on the time-window interface plane $t=t_{i}$, which is the shared boundary of neighboring windows. For this purpose, the current window model $u_{\theta}(x,y,t)=[E_{z},H_{x},H_{y}]$ is compared with the previous window predictions at the same points, $u_{\text{prev}}(x,y,t)$, and the mean squared error is minimized:
@@
The previous window predictions are treated as fixed targets by applying ‘detach‘ to prevent backpropagation from flowing into the past; in this way, the constraint affects only the parameters of the current window. Since the initial condition in the first window is provided using an analytic reference, the interface loss is not applied there. This term is defined on a single time slice on the shared boundary plane $t=t_{i}$ of successive windows; therefore, it does not include causality-aware weighting.
Ultimately, the interface continuity loss is added to the total loss function obtained in the theoretical sections with a fixed coefficient, contributing to a physically consistent and uninterrupted temporal evolution across windows:
$\mathcal{L}=w_{\text{pde}}\mathcal{L}_{\text{PDE}}+w_{\text{bc}}\mathcal{L}_{\text{BC}}+w_{\text{ic}}\mathcal{L}_{\text{IC}}+w_{\text{int}}\mathcal{L}_{\text{INT}}$ (13)
@@ 2.3.2 Poynting Loss
At the risk of repetition, it should be stated once more that Maxwell’s equations inherently include energy conservation, causality, and finite propagation speed, yet a neural network does not possess the mathematical structure that channels this content by default. For the model, the solution space is a mathematical optimization problem, and the laws of physics are not implicitly defined within this space. If the model is not forced to preserve energy as an explicit constraint, the optimizer focuses only on minimizing the $L^{2}$ loss value. Although time-windowed training slows down the tendency to converge toward a trivial solution, it does not prevent the cumulative energy drift behavior. For this reason, during implementation, Poynting’s theorem is incorporated into the loss function as a regularization term $L_{\text{pyt}}$.
A Poynting-theorem-based loss implementation can be carried out using two different approaches: global and local. Both approaches rely on the same physical quantities, but they integrate conservation information into optimization in different ways. The global approach produces a single scalar balance from an integral defined over the domain, whereas the local approach enforces the same balance as a residual at each collocation point. Under the non-dimensionalization scheme of Appendix B, the characteristic magnetic field scale $H_{0}=E_{0}/Z_{0}$ ensures that the electric and magnetic energy contributions are equally weighted. Consequently, the non-dimensional energy density simplifies to:
@@ 3.2 Results
Multiple versions of the PINN model developed in this study were trained under two primary scenarios and with different implementation approaches. The main version from which the other variants are derived will be referred to as pinnA.
One of the most fundamental challenges faced by time-marching PINN models in the literature is the error accumulation problem. [CHEN2024111423] In the continuous lossless PEC setting, total energy is constant; therefore deviations in discrete solvers are treated as numerical error (with FDTD providing the reference baseline in this study). In neural network applications for Electromagnetics, this error is often characterized as energy drift, meaning a gradual attenuation of energy over time. A key reason is the tendency of neural networks to converge toward trivial solutions. Rather than amplifying field magnitudes, the network often suppresses them, leading to amplitude damping and, consequently, energy loss. [CHANG2024112791]
While the time-marching strategy is effective at preventing causality collapse, it can produce a cumulative drift pattern. This should not be seen as an anomaly, but rather as a natural consequence of how the method operates. In time marching, the terminal values of each time window are treated as the initial conditions for the next window. As a result, even small errors formed in one stage are inherited by the subsequent window. Energy loss therefore becomes cumulative and deepens as time progresses.
@@
In the first part of this section, the configuration of pinnA is presented and the accuracy and energy metrics are summarized. In the second part, the techniques used to resolve the cumulative energy drift problem are described in detail. In the final part, we discuss the lossy variant pinnL, the four-window variant pinnW, and pinnP, in which the phenomenon hereafter referred to as the parenthesis effect is observed.
The variants are named according to their distinguishing feature: pinnA (baseline with all proposed techniques), pinnL (lossy medium), pinnW (four-window configuration), and pinnP (parenthesis effect analysis).
3.2.1 pinnA

### v2

@@ Introduction
Maxwell’s equations are a four-part system of differential equations that forms the cornerstone of classical electromagnetism and describes the space-time dynamics of macroscopic electromagnetic fields. This system of equations reveals the fundamental relationships among the electric field vector E, the magnetic field vector H, the electric flux density D, the magnetic flux density B and the current density J. The electromagnetic properties of the medium are defined by the parameters electric permittivity ($\varepsilon$), magnetic permeability ($\mu$) and electrical conductivity ($\sigma$). The differential-form expressions of the equations are as follows.
@@
Different numerical techniques such as the finite difference time domain (FDTD) [19], finite element method (FEM)[4] and method of moments (MoM) [6] have been developed for modeling and analyzing electromagnetic fields. These conventional approaches are based on the discretization of Maxwell’s equations and iterative solution algorithms. Therefore, they require intensive computational resources in problems involving complex geometries and variable material properties.
Physics-Informed Neural Networks (PINNs) were introduced between 2017 and 2019 by Maziar Raissi, Paris Perdikaris and George E. Karniadakis [16]. The PINN method involves directly incorporating physical laws defined by partial differential equations into the neural network training process. Although it can be formulated within a supervised learning framework, it generally reduces dependence on labeled data since it is performed through optimization residuals.
In the literature, adaptations of PINN applications to the time-dependent Maxwell equations have remained limited due to issues such as spectral bias, the sensitivity of electromagnetic problems to scaling, causality violations and loss balancing. The aim of this study is to present a hybrid methodology that addresses the common issues observed in PINN applications for the time-dependent Maxwell equations.
In section 1, a well-posed initial-boundary value problem formulation is obtained for the TMz mode of the time-dependent Maxwell equations. The non-dimensionalization steps are carried out and the main physics-informed loss components used for the resulting coupled PDE system are defined.
@@
In section 3, the metrics used in the study are introduced; the performance of the PINN model is evaluated through field-based error metrics and energy metrics. In addition, three different variants of the Poynting regularizer developed in this study are examined under the ablation studies section.
@@ 1.1 Well-Posedness
In their most general form, Maxwell’s equations form a coupled system of partial differential equations. Without appropriate initial and boundary data, the problem is underdetermined and admits infinitely many solutions. Neural networks rely on the universal approximation theorem and can essentially be viewed as function approximators that perform function-space mapping. [8]
For this reason, in order to solve Maxwell’s equations with PINN architectures, the problem that the model will learn must first be formulated in a well-posed manner. According to Hadamard’s classical approach, three criteria must be satisfied for well posedness: (1) existence of a solution, (2) uniqueness of the solution, and (3) continuous dependence of the solution on the data. [5]
In this study, the 2D time-domain Maxwell equations are considered in the $TM_{z}$ mode. A rectangular cavity with PEC walls is used, and a Gaussian pulse initialization is applied to excite the modal content. Therefore, under the source-free and lossless assumption ($J=0,\sigma=0$), and together with appropriate initial and boundary conditions, the resulting system defines a well-posed initial–boundary value problem:
@@
Modal decomposition is a common strategy in PINN formulations for electromagnetic problems, as it facilitates mode-specific learning and improves numerical conditioning. [15] The process of obtaining the well-posed system of equations and the modal decomposition step are described in detail in Appendix A.
@@ 1.2 Physics-Informed Loss Components
For electromagnetic cavity resonator modeling, the loss function formulation consists of partial differential equation (PDE), initial condition (IC) and boundary condition (BC) components. The PDE loss component is defined through residual functions that enforce Maxwell’s curl laws. The boundary conditions loss, which varies depending on the specific scenario, enforces whether the equations satisfy the physical laws at the boundaries. The initial condition loss regularizes the transient electromagnetic response inside the cavity according to the chosen initialization type. In this section, the theoretical development of the loss components is addressed.
@@ 2 Implementation
However, in the implementation stage, this changes. When training PINNs, the infinite dimensional solution space must be approximated using finite parameters and integral norms must be optimized using sampled estimators. In this discrete, parametric and local optimization based setting, the guarantees provided by theory may be insufficient. Scenarios that contradict physical reality are entirely plausible. For example, the optimization algorithm could violate energy conservation while minimising PDE residuals; the neural network could allow information to flow backwards in time; or field discontinuities could arise at material interfaces. These scenarios do not originate from a deficiency in the theoretical formulation, but from the intrinsic limitations of approximation and sampling based learning in neural networks. Consequently, the main goal at the implementation stage is to translate the physical guarantees implicit in theory into explicit constraints that guide the optimizer.
@@ 2.1 Training Pipeline
The core of the model is a deep fully connected neural network with 8 hidden layers, each consisting of 128 neurons. During training, to improve gradient flow and to overcome the optimization difficulties of deep networks, a skip connection is applied after every second hidden layer inspired by the ResNet architecture [7]. To ensure robust performance across different geometries, a set of problem-specific engineering approaches has been adopted:
Input normalization. The tanh activation function is symmetric around zero and yields its strongest gradients in this region. As inputs move away from zero, tanh enters saturation and gradients diminish. This imbalance causes some inputs to contribute more to the gradients than others. To prevent this issue, the spatial coordinates are dynamically normalized to the range [1,1][-1,1] depending on the cavity aspect ratio $\beta=b/a$. [11] [9]
Sinusoidal time features. Electromagnetic wave propagation in a closed cavity inherently exhibits periodic behavior. To explicitly encode this structure, the temporal input is enriched with the sinusoidal features $\sin(2\pi t/T)$ and $\cos(2\pi t/T)$. This approach provides the network with a natural basis for representing oscillatory field evolution. [17]
Output scaling. During the non-dimensionalization stage, the relationship between the electric and magnetic field components was defined as $H_{0}=E_{0}/Z_{0}$, where $Z_{0}$ is the characteristic impedance of free space. When the neural network’s output layer is initially initialized with random weights, all outputs produce values of similar magnitudes. However, physically, the $E_{z}$ and $H_{x},H_{y}$ components are expected to have different scales. To eliminate this inconsistency, the magnetic field components are scaled by a factor of 0.1 in the output layer. This design facilitates the optimization process by ensuring that the network’s starting point lies in a physically meaningful region. [12]
@@
Two-stage optimization. The optimization process is based on a two-stage strategy that combines the advantages of first-order and second-order methods. In the first stage, the global structure of the solution is learned for 1500 epochs using the Adam optimizer. In this stage, gradient clipping is used to eliminate the risk of gradient explosion and an LR scheduler that dynamically reduces the learning rate on plateaus where convergence slows down is employed. In the second stage of optimization, starting from the point reached by Adam, high-precision fine-tuning is performed with the L-BFGS optimizer, which uses local curvature information of the loss landscape.
@@
Dynamic sampling. To maximize both the efficiency and the accuracy of the training process, both sampling strategies and loss function weights are managed dynamically. Collocation points, at which the PDE residual is evaluated, are adapted according to the cavity aspect ratio, and their density is increased for narrow geometries.
@@
Similarly, to preserve causality during transitions between time windows, denser sampling is performed in regions close to the beginning of each window. During the early epochs of training, the weight of the loss term representing the initial conditions $({L}_{IC})$ is increased, while in later epochs the weight of the loss term representing the PDE $({L}_{PDE})$ is gradually raised. This integrated approach enables the model to first learn the overall structure of the problem and then refine its solution to achieve full consistency with the governing physical equations.
@@ 2.2 Causality of PINNs
When the retarded solutions of Maxwell’s equations are adopted, the field value at each point depends only on the initial and boundary data within that point’s past light cone. In other words, the field values at $t=1.0$ are formed as a function of the values at $t<1.0$. This behavior is a natural consequence of causality in hyperbolic systems.
@@
Various approaches have been explored in the literature to prevent this issue. Examples include causal training [18], which directly encodes causal information flow into the loss design; time marching [13] and causal sweeping [14], which arrange training sequentially from past to future; XPINNs-type decompositions [10], which split the domain into spacetime subregions; and curriculum learning [1], which aligns optimization by gradually changing difficulty or sampling weights.
In this study, a time-domain decomposition strategy is adopted to prevent causality violation. This strategy is applied to the model through the time marching approach.
@@ 2.2.1 Time Marching
The time marching method partitions the temporal domain into sequential windows, training each window separately. This approach is inspired by the time stepping concept in FDTD methods, while preserving the mesh-free advantages of PINNs.
@@ 2.2.3 Causality-Aware Weighting
One of the important issues in PINN training is the phenomenon of forward error accumulation, where the error grows as it is propagated forward in time. When time marching is used, even if each window is solved within itself, a small PDE inconsistency at the beginning of the window can spread through derivative terms and become pronounced at mid to late times. For this reason, a causality-aware weighting mechanism is proposed in this study. This mechanism can be considered a simplified and deterministic variant of the causal training approach.
@@
In practice, with the choice $\gamma=2$, samples at the beginning of the window $\tau=0$ receive weight $1$, while those at the end $(\tau=1)$ receive approximately $e^{-2}\simeq 0.135$. This creates a weight ratio of approximately $7.4\times$ between the beginning and the end of the window. As a result, stochastic gradient estimates systematically focus on the beginning of the window, optimization suppresses early time PDE inconsistencies first, and an equation-consistent initial regime is established.
@@ 2.3.1 Interface Continuity Loss
To reliably stitch successive windows, a two-stage continuity mechanism is employed. The first stage is the IC continuity loss applied at the start of each window. In the first window, the initial-condition term enforces the physical initial state by prescribing a Gaussian pulse and zero magnetic field. In subsequent windows, the same term no longer targets an analytic profile, but instead matches the trained PINN solution from the previous window. Accordingly, for successive time windows $\mathcal{W}_{i}=[t_{i},t_{i+1}]$, the $E_{z}$, $H_{x}$, and $H_{y}$ fields at the beginning of window $k$ are penalized in terms of squared error against the fields at the terminal time of window k1k-1. The fields imported from the previous window are detached during backpropagation so that they serve as fixed targets; gradients therefore flow only through the parameters of the active window, keeping the training scheme temporally causal.
The second stage is a dedicated interface continuity loss defined on an independent sampling set at the window boundary. This term compares the current window model $u_{\theta}(x,y,t)=[E_{z},H_{x},H_{y}]$ with the previous window predictions $u_{\text{prev}}(x,y,t)$ at interface centered points $(x_{\text{int}},y_{\text{int}},t_{\text{int}})$ and the mean squared error is minimized:
@@
This enforces continuity more strongly over a narrow spatiotemporal band around the interface, rather than relying solely on a single IC slice defined at the window start. The weights $w_{\text{ic}}$ and $w_{\text{int}}$ are used to tune the relative influence of these two continuity layers within the overall loss. Since the initial condition in the first window is provided using an analytic reference, the interface loss is not applied there. Both terms are defined on a single time slice on the shared boundary plane $t=t_{i}$ of successive windows; therefore, they do not include causality-aware weighting.
@@ 2.3.2 Poynting Loss
At the risk of repetition, it should be stated once more that Maxwell’s equations encode the information of energy conservation, and through their hyperbolic structure, the principle of finite propagation speed; when physically meaningful retarded solutions are adopted, the framework of causality becomes clear. A neural network, however, does not possess this mathematical structure by default. From the network’s perspective the solution space is simply an optimization problem, and the laws of physics carry no intrinsic guarantee within it. If the requirement of energy conservation is not imposed as an explicit constraint, the optimizer in practice focuses only on minimizing the $L^{2}$ loss value. Although time-windowed training slows the tendency to converge toward a trivial solution, it cannot reliably prevent cumulative energy drift. For this reason, during implementation, Poynting’s theorem is incorporated into the loss function as a regularization term $L_{\text{pyt}}$.
A loss implementation based on Poynting theorem can be carried out using two different approaches: global and local. Both approaches rely on the same physical quantities, but they integrate conservation information into optimization in different ways. The global approach produces a single scalar balance from an integral defined over the domain, whereas the local approach enforces the same balance as a residual at each collocation point. Under the non-dimensionalization scheme of Appendix B, the characteristic magnetic field scale $H_{0}=E_{0}/Z_{0}$ ensures that the electric and magnetic energy contributions are equally weighted. Consequently, the non-dimensional energy density simplifies to:
@@ 3.2 Results
Multiple versions of the PINN model developed in this study were trained under three scenarios and with different implementation approaches. The main version from which the other versions are derived will be referred to as pinnA.
One of the most fundamental challenges faced by time marching PINN models in the literature is the error accumulation problem. [3] In the continuous lossless PEC setting, total energy is constant; therefore deviations in discrete solvers like FDTD are treated as numerical error.
In neural network applications for electromagnetic problems, this error is often characterized as energy drift and most commonly manifests as decay. The underlying reason is the tendency of neural networks to converge toward trivial solutions. Rather than amplifying field magnitudes, the network tends to suppress them; this leads to amplitude damping and energy loss. [2]
Although the time marching strategy is successful in preventing causality collapse, it causes energy drift to become cumulative. This situation can be seen as a natural consequence of how the method operates. In the time marching method, the terminal values of each time window are treated as the initial condition for the next window. This causes every small error formed in the previous stage to be inherited by the subsequent window. Therefore, energy drift becomes cumulative and deepens as time progresses.
@@
3.2.1 Main Model